Home >Documents >2011-12 winter shutdown planning A.Tauro 6-1-2012 6/1/12ALICE 2011-12 winter shutdown schedule -...

2011-12 winter shutdown planning A.Tauro 6-1-2012 6/1/12ALICE 2011-12 winter shutdown schedule -...

Date post:18-Jan-2018
Category:
View:214 times
Download:0 times
Share this document with a friend
Description:
Dec JanFebMar Wk Mo 5 After 17: After 17: Tu We Th After 17:00 Fr AUG 48V (EL) Sa DSO Su Xmas Day NY Day 17/1/12TC meeting - A.Tauro UX25 access restrictions Access restrictions: Removal/installation of shielding (restriction includes RB24 triplets) PX24 blocks: no access CR5 and UX25 PX24 beams (35t): no access CR1-5, Expo, and UX-25 All other blocks: no access to UX-25 (or 1-2 teams max, not on A-side) TRD insertion: no access to MNF and L3 AUG test (Arrêt Urgence General – General Emergency Stop) – no access UX25 DSO test (Department Safety Officer) – no access UX25 No access SX2 during manipulation of blocks in SX2 (except ACR and WRs) Close L3 Open L3 Legend No access (except transport) Access possible No access 3 To be confirmed
Transcript:

winter shutdown planning A.Tauro /1/12ALICE winter shutdown schedule - A.Tauro 1 Dec JanFeb Mar Wk Mo Tu We Th Fr Sa Su Xmas DayNY Day winter shutdown 17/1/12TC meeting - A.Tauro Close L3 Open L3 Beam off: Dec 8 th at 6:00 AM Beam on: Mar 7 th at 18:00 PM 2 Dec JanFebMar Wk Mo 5 After 17: After 17: Tu We Th After 17:00 Fr AUG 48V (EL) Sa DSO Su Xmas Day NY Day 17/1/12TC meeting - A.Tauro UX25 access restrictions Access restrictions: Removal/installation of shielding (restriction includes RB24 triplets) PX24 blocks: no access CR5 and UX25 PX24 beams (35t): no access CR1-5, Expo, and UX-25 All other blocks: no access to UX-25 (or 1-2 teams max, not on A-side) TRD insertion: no access to MNF and L3 AUG test (Arrt Urgence General General Emergency Stop) no access UX25 DSO test (Department Safety Officer) no access UX25 No access SX2 during manipulation of blocks in SX2 (except ACR and WRs) Close L3 Open L3 Legend No access (except transport) Access possible No access 3 To be confirmed Dec JanFebMar Wk Mo 5 After 17: No ventilation After 17: Tu We Th After 17:00 Fr No ventilation AUG 48V (EL) Sa DSO Su Xmas Day NY Day 17/1/12TC meeting - A.Tauro Wks and 7: access in normal hours (e.g. TPC) to be coordinated with TC Wks 1-2: no access inside L3 during TRD SMs insertion When moving from low-beta to insertion platform: no access to MNF When inserting the module: no access L3 and MNF Watcher needed in wks and from wk 6 onwards L3 and MNF access restrictions Close L3 Open L3 4 To be confirmed EN-HE 8 th Dec (9:30): reinstall crane (PR-774) electronics 8 th Dec (AM): load test cavern cranes (20t) 30 th Jan. UX25 crane inspection by Brunnhuber Wk 8: test UJ22 jig + annual maintenance P2 cranes 6/1/12ALICE winter shutdown schedule - A.Tauro EN-HE-LM 14 th Dec: maintenance lift CR1-4 Lift stopped whole day (7:00 16:00) Cavern lift has to be stopped as well for one day. Date tbc 5 TE-VSC 8 th and 9 th Dec: close MNF valve and venting central beampipe Feb 1 st 23 rd : pump down central beampipe 6/1/12ALICE winter shutdown schedule - A.Tauro 6 GS-ASE Week 52: evacuation test Jan. 5 th 13 th : RAMSES firmware upgrade Jan. 9 th : red phone maintenance Jan. 20 th and 26 th : test fire detection Week 5: sniffer maintenance Week 6: maintenance gas detection (UX, CR5) Feb. 21 st : maintenance PAD/MAD Guard will be provided by GS in order not to interfere with cavern access 6/1/12ALICE winter shutdown schedule - A.Tauro 7 EN-CV-OP Dec 7 th Dec at 18:00-22:00: adjustment chilled water flow in the cavern + lower L3-Dipole cooling settings Will affect all cooling plants Keep as many detectors on till the 8 th (DCS shutdown), in order to check that regulation works (EN-CV-DC side) Keep L3 and Dipole till the end of the test Week 50: rinsing of mixed water in UX25 cavern (Philippe) Refilling in parallel will not affect ALICE cooling End wk 50 (16 th Dec): test of Dipole flooding alarm Need access to Dipole (Philippe) 6/1/12ALICE winter shutdown schedule - A.Tauro 8 EN-CV-OP Jan/Feb Tue 2 nd and Wed 3 rd Jan: Maintenance ITS (old) ventilator in CR5 Thu 5 th and Fri 6 th Jan: maintenance PMD-MTG-MTK ventilation units Fri 6 th Jan: stop chilled water (SU2) Therefore stop mixed water (cavern, CRs) Switch tap water into CRs Week 2: Cleaning cooling towers (Mon to Thu) Maintenance L3 injection and UX25 extraction/injection units. Test of alarms New motor L3 ventilation (date tbc) Modification of chilled water pipe for L3 jacket (Mon) Maintenance mixed water pumps PX24 (Mon-Wed) and UW25 (Wed-Thu) Preparation services new ITS dehumidifier Consolidation of automatic switchover to tap water into CRs Week 3: Restart chilled & mixed water (around Wed, tbc) Maintenance L3 extraction unit. Test of alarms Maintenance ventilation unit SG2 (Mon-Tue) will cut Ar/H2 6/1/12ALICE winter shutdown schedule - A.Tauro 9 EN-CV-OP Repair leak FB pipe: 23 rd Jan Fernando compensatory measures Removal unused FB pipe (I-side): 24 th -25 th Jan Need nacelle 6/1/12ALICE winter shutdown schedule - A.Tauro 10 EN-EL 1/2 13/12: Test general secours: stop general services for 10 Will affect all general services, also at P2. The UPSes for DAQ and DCS/HLT will bridge the power cut (assuming they manage to repair the DCS/HLT one on ). Also rack power (CRs and UX) will not be affected. However all 'wall outlets', WR of the ACR, offices, meeting room, SXL2, coffee machine will be affected... Also the non ups part of the cooling plants as well as the gas system will be affected 17/12: Test auto-transfert: 1.5 days to configure + 1 day test No impact on experiments Wed 18 th Jan: maintenance power transformers: Will cut EXD1/2X for 3h cut power to CRs (normal network) Will cut ITS ventilator and HMPID rack in CR5 Test AUG P2: Fri Jan 20 th No access UX25 Major electrical perturbation affecting P2 for whole day 6/1/12ALICE winter shutdown schedule - A.Tauro 11 EN-EL 2/2 Consolidation 48V network (after incident 11 th Jan 2011) 1 day no access to UX25 Feb 17 th (tbc, K.Foraz) Maintenance power transformer EMT306/2X Feb 17 th 6:00 to 10:00 Will stop UX25 air extraction and injection (4h) No access UX25 from 6:00 to 10:00 6/1/12ALICE winter shutdown schedule - A.Tauro 12 EN-CV-BE New ITS dehumidifier unit in SX2 Electrical control cubicle during week 2 Power cubicle : delivery week 4 Cabling : week 4/5 Start-up and commissioning : week 6 /7 6/1/12ALICE winter shutdown schedule - A.Tauro 13 EN-CV-DC RP sampling: Dec 8 th TRD. Periodic maintenance + replacement chilled water valve + preventive replacement pump + cleaning vacuum pump: plant stopped 2 days on Jan 5 th and 6 th SPD. Periodic maintenance + put entire plant on UPS: plant stopped 1 day on Jan 6 th TOF-PHOS-EMC-SPD. Installation PHOS circulator + replacement flowmeter + periodic maintenance: plant stopped from 9 th to 17 th Jan TPC. Periodic maintenance + preventive replacement pump + include heaters in supervision + cleaning vacuum pump: pant stopped from 18 th to 20 th Jan HMPID. Periodic maintenance + preventive replacement pump: plant stopped 2 days on Jan 23 rd and 24 th TPC RR. Periodic maintenance + install new filters: plant stopped from Jan 25 th to Feb 2 nd SDD-SSD. Periodic maintenance + replacement pump: plant stopped from 3 rd to 9 th Feb 6/1/12ALICE winter shutdown schedule - A.Tauro 14 BE-ABP-SU (survey) Week 50 (Sat-Sun): with L3 doors open: Zero measurement (SF, TPC feet, MNF valve, MNF) Low Beta network update C-side photogrammetry (EMCal) Weeks 1-2: after each TRD insertion, requiring day on low-beta: MNF valve and MNF Week 4: survey MTG Week 5: at the end of TRD and EMCal, requiring three full days with both doors open: TRD modules, CAL frame photogrammetry, MNF valve, MNF, SF deformation, TPC feet Week 6: V0-FMD1 (Mon) and MTK ST1-2 (Wed-Fri) Week 7: no access (reserve) Weeks 9-10:survey MTK ST3-4-5 and PMD 6/1/12ALICE winter shutdown schedule - A.Tauro 15 DAQ DAQ shutdown 21 st Dec mid-day DAQ start after cooling and power interventions, end wk 3 Arrangements for individual LDCs can be organized (send a mail to 6/1/12ALICE winter shutdown schedule - A.Tauro 16 DCS Week 49 Dec. 8 morning: shutdown all DCS for UPS repair; restart a.s.a.p. (afternoon) Profit from DCS stop to make backup ! Start (hardware) installation of new DCS cluster Week 50 Dec. 1214: Operating system and PVSS upgrade Reboots required, coordinated with detectors Dec. 15: New framework available for detectors Detectors to validate functionality with new software Week Preparation of new DCS cluster DCS available during CERN End-of-Year closure with usual access limitations and best effort support Week New DCS cluster available for detectors Expect all detectors to run on new cluster by week 5 6/1/12ALICE winter shutdown schedule - A.Tauro 17 TPC Take out capacitors on all A/C ROC 6/1/12ALICE winter shutdown schedule - A.Tauro 18 TPC heat screen Check the loops for the thermal screen in week 6 (there was a leak) We will need: access to the TRD cooling plant access to C side and A side of TPC access to the miniframe and PP0 Since the pipes are on the outside of the miniframe we would probably need a nacelle. 6/1/12ALICE winter shutdown schedule - A.Tauro 19 TPC sectors 11 to 15 Week 49: Thu 8/12: displace FMD2 box. The box can be reinstalled anytime (at latest in week 5) TPC A-14/15 (4 days anytime) Week 3: move PMD to parking position (see next slide) Week 4: Mon 23/1 AM: disconnect and displace SPD box Mon 23/1 PM: disconnect and displace SSD box Tue-Fri: TPC A-11/12 (4 days) Sat: put back and test SSD box Week 5: Mon 30/1: put back and test SPD box Tue: move V0-FMD1 to parking position (PMD already in parking position) Wed-Fri: TPC A-13 (2 and 1/2 days) + FMD2 survey Sat: move back and test V0-FMD1 Week 6: Mon 6/2: survey of V0-FMD1 6/1/12ALICE winter shutdown schedule - A.Tauro 20 PMD Thu 19 th Jan: PMD moves to Parking Position Week 4 - 6: PMD Maintenance Replacing some of the modules which are having problem Replacing some of the FEE Chains Need cooling for the DAQ (Crocus) so that we can check the health of the modules and FEE chains. Week 7 (after 17:00) and 8: Access to Crocus Crate for DAQ Week 9 -10: PMD Moves back to Final Position 6/1/12ALICE winter shutdown schedule - A.Tauro 21 T0 Investigation of the problem with laser system (laser, attenuator, optical splitters, optical cables) Rack 018,019, start from ~16 January Miniframe (from Feb.7 th ) Upgrade of electronics and check of spare modules Rack 018,019, start from ~16 January 6/1/12ALICE winter shutdown schedule - A.Tauro 22 TPC laser Annual laser maintenance, 1-2 days. Needs normal power and cooling Open mirror box below TPC sector A13 (bottom below TPC endplate) and possibly replace camera or mirror. 1-2 days. I want to do this because a weak picture from this mirror/camera makes C- side alignment difficult. Best done with L3 doors open. I believe access between the TOF and TPC should be OK. I need to be able to run the laser, so we need to mount t

Click here to load reader

Reader Image
Embed Size (px)
Recommended